OpenStudy (anonymous):

-1/6(x-12)+1/3(x+3)=x+18

OpenStudy (anonymous):

-1/6x+2 + 1/3x+1=x+18 -1/6x+1/3x-x=18-2-1 -5/6x=15 x=-18